# SplitVAEs: Decentralized scenario generation from siloed data for stochastic optimization problems

### v1

@@ VI Experimental Results
Our experimental results are primarily geared towards establishing the computational efficiency of the SplitVAEs method while delivering scenarios of the same statistical quality as the underlying dataset. In this section, we demonstrate the efficacy of the SplitVAEs approach on multiple open-source datasets using large-scale distributed memory-based experiments. All experiments were conducted using OpenMPI [33] and mpi4py [34] for distributed processing while leveraging PyTorch [35] for constructing, training and evaluating the machine learning models.
VI-A Applicability to a Diversity of Use Cases
To demonstrate the broad applicability of our proposed framework, we primarily evaluate the decentralized scenario generation capability of our proposed framework using the USAID [36], ACES [37] and ACTIVSg datasets [38]. The USAID dataset comprises of time series data of health commodity supplies for different countries. From the public ACES dataset we extract time series data for ($\text{CO}_{2}$) emissions from 25 refineries in the US Gulf Coast for Summer 2017. Lastly, the ACTIVSg dataset provides the Demand and Renewable generation dataset for a power transmission network with an overall footprint across Texas. We focus on hourly demand of 1125 individual network buses and renewable production from 87 renewable generators respectively which are referred to as Demand and Renewable datasets respectively.
TABLE I: Means (with standard deviations measured in $1e^{-3}$) of various evaluation scores across four different datasets.
@@ VI-A Applicability to a Diversity of Use Cases
Table I presents the mean and standard deviation of each evaluation metric for 100 iterations with dimension of edge level autoencoders set to 20. It can be observed that the performance of the SplitVAEs approach is nearly identical to that of centralized training schemes. As a result, we can expect our method to generate high quality scenarios for different application areas including but not limited to supply chains, carbon credit compliance planning and forecasting renewable energy production for power systems.
@@
Additionally, Figures LABEL:fig:t-SNE, LABEL:fig:tsa_centroid, and LABEL:fig:tsa_autocorrelation present a statistical perspective to compare the observed data with scenarios generated by three different methods across four case studies. Figure LABEL:fig:t-SNE illustrates density curves, estimated using a Gaussian-based kernel method, of one-dimensional embeddings extracted from generated scenarios and observed data by the t-SNE method [26]. Meanwhile, Figures LABEL:fig:tsa_centroid and LABEL:fig:tsa_autocorrelation provides a comparison of centroids and autocorrelation coefficients, respectively. According to Figure LABEL:fig:t-SNE, all methods capture the underlying distributions well, with 95% of the principal components falling within the same 95% confidence range. Some minor misalignments, likely due to feature compression in the latent dimension, do not significantly affect the overall alignment between the generated scenarios and the observed data. Furthermore, Figures LABEL:fig:tsa_centroid and LABEL:fig:tsa_autocorrelation, show a close alignment between scenarios generated by SplitVAEs and state-of-the-art benchmark methods while capturing interdependencies present in original datasets both in terms of centralized and autocorrelations. To illustrate a practical realization, we present seasonal carbon emission scenarios from a refinery at Port Arthur, Texas in Figure LABEL:fig:PortArthur. As depicted in Figure LABEL:fig:PortArthur, the spatiotemporal information from scenarios generated by our SplitVAE closely aligns with observed data and other benchmark methods.
VI-B Reduction in Data Movement
@@
To evaluate the transmission efficiency of SplitVAEs, we measure the total payload size handled by the system after a full cycle of our Algorithm 7 (forward propagation from edge-level encoders to the server-level VAE, and then from the server-level VAE to edge-level decoders; backpropagation from edge decoders to the server-level VAE, and finally from the server-level VAE to edge encoders). Figure 6 compares the original dataset size with the transmission overhead for various edge dimension choices. The results show a significant reduction in data size as latent dimensions decrease. For instance, in the Demand dataset, the transmission size drops from 1830 MB for the original data to 475 MB at latent size 20, 223 MB at size 16, and 154 MB at size 8. Reductions at the smallest dimensions are 4.7x, 2.3x, 2.8x, and 2.7x for the USAID, ACES, Demand, and Renewable datasets, respectively. Combined with Figure LABEL:fig:t-SNE and Table I, these findings show that SplitVAEs deliver high quality scenarios with up to 5x reduction in transmission size.
VI-C Handling Dimensional Heterogeneity
@@
In this case study, we evaluate the ability of SplitVAEs to handle heterogeneous data dimensions present across various stakeholders. Dimensional heterogeneity refers to differences in the size and structure of datasets across various edge devices in a decentralized network. For our experiments, we decompose the Demand dataset into regions with varying numbers of power system buses, where each bus represents a unique temporal demand curve. This results in multiple region decomposition cases, with each region acting as a distinct edge device handling a different dimensional dataset. We then conduct experiments using three specific region decompositions of the Demand dataset: 20 regions, 60 regions, and 120 regions. Each configuration represents a different level of dimensional heterogeneity, with a varying number of buses (or demand curves) per region.
@@
Figure LABEL:fig:decomposition_spatiotemporal provides a visual representation of the data, depicted as means (solid circle) and standard deviations (whiskers), generated by the SplitVAEs model across the three different regional decompositions. The plots demonstrate that the model can adapt to varying local dataset dimensions, effectively capturing the underlying data distributions in all cases. Therefore, the results suggest that the SplitVAEs framework is robust to dimensional heterogeneity, delivers high performance across different regional decompositions.
For this case study, we present Figure 9 that illustrates the global training loss for the SplitVAEs model across different regional decompositions of the Demand dataset. The x-axis represents the number of global training epochs, while the y-axis shows the total training loss, which includes both reconstruction loss and KL divergence loss. Each line in the graph corresponds to a different regional decomposition setting. As observed, all trends show initial losses that decrease over subsequent epochs, indicating the model’s convergence.
VI-D Architectural Robustness
@@ VII Conclusion
In the paper we present SplitVAEs, a novel decentralized method to generate spatiotemporally interdependent scenarios from siloed data specifically geared toward solving SO problems. Using backpropagation, we effectively decompose the training across edge-level autoencoders and a server-level variational autoencoder and enables the bi-directional insights flow among the edge and server models. We demonstrate the applicability of SplitVAEs specifically in the context of multi-stakeholder infrastructure systems by leveraging real-world datasets from various areas including supply chains, energy, and environmental science. We show that SplitVAEs generate scenarios that closely align with the real underlying data distributions by benchmarking with state of the art centralized methods and numerous statistical metrics. Using large scale, distributed memory driven experiments, our results indicate that SplitVAEs can significantly reduce data transmission, scale efficiently and help break down data silos. Overall, our analysis demonstrates that SplitVAEs are a compelling alternative that enables scenario generation for SO problems in multi-stakeholder-led infrastructure systems without the need to transfer underlying datasets.

### v2

@@ VI Experimental Results
Our experimental results are primarily geared towards establishing the computational efficiency of the SplitVAEs method while delivering scenarios of the same statistical quality as the underlying dataset. In this section, we demonstrate that SplitVAEs deliver scenarios possessing similar statistical qualities to the observed data, while showcasing their ability of handling heterogeneous datasets in a scalable fashion.
VI-A Hardware and Software Environment
All experiments were conducted using the Pete High Performance Computing infrastructure from Oklahoma State University (OSU). All experiments pertaining to large-scale distributed processing were performed using the OpenMPI library [34] coupled with Python bindings facilitated by mpi4py [35]. We utilize the PyTorch library [36] for constructing, training and evaluating the DNN models, and the inbuilt Ray Tune framework [37] for hyperparameter tuning. Specifically, we first tune the hyperparameters of the Centralized-VAE, then adapt these hyperparameters for different numbers of edge locations and desired latent dimensions in SplitVAEs.
VI-B Description of Key Datasets
[…]
VI-C Applicability to a Diversity of Use Cases
Table I presents the mean and standard deviation of each evaluation metric for 100 iterations, with output dimension of edge level autoencoders set to 20. The experiments for each dataset comprises varying number of edge nodes ranging from 25 for ACES to 1125 distinct edge processes in the Demand dataset. From Table I, we can clearly observe that the performance of our proposed framework is nearly identical to those of centralized training schemes. Thus, we can expect SplitVAEs to generate high-fidelity scenarios for different application areas including but not limited to supply chains, carbon credit compliance planning and forecasting renewable energy production for power systems.
Table I: Means (with standard deviations measured in $10^{-3}$) of various evaluation scores across four different datasets.
@@ VI-C Applicability to a Diversity of Use Cases
Additionally, Figures LABEL:fig:t-SNE, LABEL:fig:tsa_centroid, and LABEL:fig:tsa_autocorrelation present a comparison of the observed data with scenarios generated by three different methods across four case studies. Figure LABEL:fig:t-SNE illustrates density curves, estimated using a Gaussian-based kernel method, of one-dimensional embeddings constructed from generated scenarios and observed data by the t-SNE method [27]. Meanwhile, Figures LABEL:fig:tsa_centroid and LABEL:fig:tsa_autocorrelation provides a comparison of centroids - computed by averaging the quantity of interest for all the nodes at each distinct time point - and their corresponding autocorrelation coefficients, respectively.
@@
From Figure LABEL:fig:t-SNE, we can observe that all methods capture the underlying distributions well, with 95% of their principal components falling within the confidence range. Some minor misalignments, primarily result from feature compression in the latent dimension, do not significantly affect the overall alignment between the generated scenarios and the observed data. Further, Figures LABEL:fig:tsa_centroid and LABEL:fig:tsa_autocorrelation, show a close alignment between scenarios generated by SplitVAEs and state-of-the-art benchmark methods while capturing interdependencies present in original datasets both in terms of centralized and autocorrelations.
To illustrate a practical application, we present seasonal carbon emission scenarios from a refinery at Port Arthur, Texas, in Figure LABEL:fig:PortArthur. As shown, the scenarios generated by the SplitVAEs closely align with the observed data and other benchmark methods. Specifically, within the perimeter of the industrial complex, the observed emission profile is approximately 180 to 190 Megagrams (Mg), which is accurately reflected in the scenarios generated by all three methods. Moreover, the emission patterns derived from the scenarios generated by the SplitVAEs follow the same trend as the observed data.
VI-D Reduction in Data Movement
@@
To evaluate the communication or transmission efficiency of SplitVAEs, we measure the total payload size handled by the system after a full cycle of our Algorithm 7 comprising one forward pass of embeddings followed by backward pass of gradients and errors. Figure 6 compares the original dataset size with the transmission overhead for various edge dimension choices. The results show a significant reduction in data size as latent dimensions decrease. For instance, in the Demand dataset, the transmission size drops from 1830 MB for the original data to 475 MB at latent size 20, 223 MB at size 16, and 154 MB at size 8. Data transmission reduction factors at the smallest dimensions are $4.7$, $2.3$, $2.8$, and $2.7$ for the USAID, ACES, Demand, and Renewable datasets, respectively with respect to original data sizes. Combined with Figure LABEL:fig:t-SNE, Figure LABEL:fig:tsa_autocorrelation, and Table I, these findings demonstrate that SplitVAEs deliver high-fidelity scenarios with up to nearly a 5 times reduction in data transmission costs.
VI-E Handling Dimensional Heterogeneity
@@
In this case study, we evaluate the ability of SplitVAEs to handle heterogeneous data dimensions present across various stakeholders. Dimensional heterogeneity refers to differences in the size and structure of datasets across various edge devices in a decentralized network. For our experiments, we decompose the Demand dataset into regions with varying numbers of power system buses, where each bus represents a unique temporal demand curve. This results in multiple region decomposition cases, with each region acting as a distinct edge device handling a different dimensional dataset. Our decomposition cases rely on graph-partitioning schemes [41] that attempt to decompose the power network graph by minimizing the total number of inter-regional transmission lines. As a result, we obtain three different decompositions pertaining to $n=20$ regions, $n=60$ regions, and $n=120$ regions wherein, each region represents a distinct SplitVAE edge node and the dimension of local data at each node is non-homogeneous.
@@
Figure LABEL:fig:decomposition_spatiotemporal provides a visual representation of the data, depicted as means (solid circle) and standard deviations (whiskers), generated by the SplitVAEs model across the three different regional decompositions. The plots demonstrate that the model can adapt to varying local dataset dimensions, effectively capturing the underlying data distributions in all cases. Therefore, the results suggest that the SplitVAEs framework is robust to dimensional heterogeneity, delivers high performance across different regional decompositions. On the other hand, Figure 9 illustrates the global training loss for the SplitVAEs model across different regional decompositions of the Demand dataset. The x-axis represents the number of global training epochs, while the y-axis shows the total training loss, which includes both reconstruction loss and KL divergence loss. Each line in the graph corresponds to a different regional decomposition setting. As observed, all trends show initial losses that decrease over subsequent epochs, indicating the model’s convergence.
VI-F Architectural Robustness
@@ VII Conclusion
In the paper we present SplitVAEs, a novel decentralized method to generate spatiotemporally interdependent scenarios from siloed data specifically geared toward solving SO problems. Using backpropagation, we effectively decompose the training across edge-level autoencoders and a server-level variational autoencoder and enable the bi-directional flow of insights among the edge and server models. We demonstrate the applicability of SplitVAEs specifically in the context of multi-stakeholder infrastructure systems by leveraging real-world datasets from various areas including supply chains, energy, and environmental science. We show that SplitVAEs generate scenarios that closely align with the real underlying data distributions by benchmarking with the state-of-the-art centralized methods and numerous statistical metrics. Using large scale, distributed memory driven experiments, our results indicate that SplitVAEs can significantly reduce data transmission, scale efficiently with growing number of edge locations and help break down data silos. Overall, our analysis demonstrates that SplitVAEs are a compelling alternative that enables scenario generation for SO problems in multi-stakeholder-led infrastructure systems without the need to transfer underlying datasets.